22. A can of Pepsi is about 11% by mass sugar (C 22 molar mass 342.3 g mol 1). Assuming the solvent is water and ignoring any other solutes, what is the mole percent of sugar in soda? (density of the soda 1.1 g mL A. 11% B. 0.65% C. 10% D. 12% E. 0.03%
Answer #1

Ans 22 B. 0.65%

Let us assume the mass of pepsi can be 100 g

so sugar in it would be 11 g

and water will be 89 g

number of moles of sugar will be 11 / 342.3 = 0.0321 moles

number of moles of water will be 89 / 18 = 4.944 moles

total number of moles = 0.0321 + 4.944 = 4.976

Mole percent of sugar = (0.0321 / 4.976) x 100

Mole percent of sugar in soda = 0.65%